# Active nematic response to a deformable body or boundary: elastic deformations and anchoring-induced flow

### v1

@@ Abstract
A body immersed in a nematic liquid crystal disturbs the fluid’s preferred molecular configuration and increases its stored elastic energy. In an active nematic, the fluid components also generate a stress in the bulk fluid. By introducing either an immersed body or boundary, a fluid flow can be triggered due to anchoring boundary conditions. The fluid imposes elastic, viscous, and active stresses on such surfaces, which, if compliant, may result in a surface deformation. We study the deformations and stress in a linearly elastic body placed in an active nematic in two-dimensions. Using complex variables techniques, exact expressions for the fluid flow, director field, surface tractions, and body deformation are derived. Qualitative differences between elastic and active stress-driven deformations are identified, depending on an active Ericksen number, anchoring conditions, and body material properties, thereby suggesting a new method for measuring the active stresses in active anisotropic biological systems. Flow profiles, external confinement, and anchoring-induced stirring are also addressed.
@@
The eukaryotic cell is host to a wide variety of deformable organelles, from the nucleus to mitochondria [1, 2, 3, 4, 5, 6]. These organelles exist in a complex, potentially anisotropic, viscous environment [7, 8], and experience mechanical stresses due to actively driven cytoplasmic flows [9]. The cell membrane itself is also deformable, and is commonly remodelled by internal processes. For instance, active anisotropic stresses are of great functional importance in cytoskeletal remodeling of the nucleus and chromatin configurations [10, 11], and in the metaphase spindle [12, 13, 14]. On a larger length scale, active systems like bacterial swarms and biofilms also exhibit anisotropy, among other complex rheological properties [15, 16, 17, 18]. Understanding how such bodies affect, and are affected by, this complex environment is of generic importance.
@@
Even when free of bodies, actively-driven anisotropic flows exhibit very rich and complex dynamics [19, 20, 21, 22, 23]. Stability [24, 21, 25, 26, 27, 28, 29, 30], topology [31], pattern formation [32, 33, 34], and mixing [35, 36] have seen substantial theoretical treatment. Confinement introduces an additional length scale, which can affect the emergent spatial structure [37, 38]. Strong active nematic stresses can also change the shape of deformable confining surfaces [39, 40, 41, 42, 43, 44, 45, 46], and they play a prominent role in directing the motility of cells [47, 48] and active droplets [49, 50, 51, 52].
@@
When an active nematic contains an immersed body, or inclusion, additional elastic stresses are immediately introduced. In passive systems, these stresses result in intricate and beautiful textures [53, 54, 55], and can be used to measure cell material properties [56, 57]. In the presence of activity, the combination of fluid elasticity and anisotropy redirects actively driven flows, leading to complex immersed body dynamics. Recent examples include an investigation of the biased rotation of gear-like bodies [58, 59], and transitions from fixed point to limit cycle to chaotic dynamics of a circular disk [60]. How body deformability affects, and is affected, by such an active environment has not been well explored.
In this letter, we present a model for an active anisotropic fluid with a soft internal or external boundary, confined to two-dimensions (2D). The model integrates Ericksen–Leslie theory [61, 62, 63] for active nematic liquid crystals (LCs), a common model for active anisotropic fluid [64, 23], and linear elasticity for deformable solids. Complex variable techniques yield novel analytical representations for the director field, velocity field, and boundary deformation. We demonstrate this for both an immersed and bounding cylinder, highlighting the insights that these analytical solutions provide.
@@
Active nematic LCs can be described by their locally-averaged molecular orientation, $\bm{n}(\bm{x},t)$, and fluid velocity $\bm{u}(\bm{x},t)$, with spatial position $\bm{x}$, time $t$, and $|\bm{n}|=1$. The director field, $\bm{n}$, evolves due to gradients in the velocity field, with strain rate tensor $\bm{\mathrm{E}}=(\nabla\bm{u}+\nabla\bm{u}^{T})/2$ and vorticity tensor $\bm{\Omega}=(\nabla\bm{u}-\nabla\bm{u}^{T})/2$, and a relaxation towards equilibrium under the action of the molecular field, $\bm{h}$. In the one-constant approximation, $\bm{h}=K\nabla^{2}\bm{n}$, for the single (Frank) elastic constant $K$. Altogether,
$\partial_{t}\bm{n}+\bm{u}\cdot\nabla\bm{n}=\bm{n}\cdot\bm{\Omega}+\lambda\bm{n}\cdot\bm{\mathrm{E}}\cdot(\bm{\mathrm{I}}-\bm{n}\bm{n})+\tilde{\bm{h}}/\gamma,$ (1)
@@
with $\gamma$ rotational viscosity, $\lambda$ tumbling/reactive parameter, and $\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\bm{h}$ the transverse part of the molecular field, where $\bm{n}\bm{n}$ is a dyadic product [61, 62, 63]. Assuming incompressibility, the fluid velocity satisfies the momentum and mass conservation equations
@@
for the pressure $p$ and elastic, viscous, and active stress tensors
$\bm{\sigma}^{e}$ $=-K\nabla\bm{n}\cdot\nabla\bm{n}^{T}+\frac{1}{2}(1-\lambda)\bm{n}\bm{\tilde{h}}-\frac{1}{2}(1+\lambda)\bm{\tilde{h}}\bm{n},$ (3a)
@@
respectively. Here, $\mu$ is the solvent viscosity, $\mu_{1}$ and $\mu_{2}$ are anistropic viscosities, and $\alpha$ is the activity strength, which can describe both extensile ($\alpha<0$) and contractile ($\alpha>0$) activity [64, 65, 23].
We non-dimensionalize the system by using an intrinsic length scale $L$ (e.g. the size of an immmersed body), velocity scale $U$ (to be defined), and stress scale $K/L^{2}$. The dimensionless system is then described by the dimensionless Ericksen number, activity strength, and viscosity ratios [66]:
@@
For small rotational Ericksen numbers, $\gamma^{\prime}\mathrm{Er}\ll 1$, the fluid flow does not affect the director field and the molecular field relaxes to equilibrium. The director evolution equation (1) takes the dimensionless form
$\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\nabla^{2}\bm{n}=\mathcal{O}(\gamma^{\prime}\mathrm{Er}).$ (5)
By further assuming small viscosity ratios, $\gamma^{\prime},\mu_{i}^{\prime}\ll 1$, the momentum equation (2a) takes the dimensionless form
$\nabla^{2}\bm{u}-\nabla p=\frac{1}{\mathrm{Er}}\nabla\cdot\bigl{(}\nabla\bm{n}\cdot\nabla\bm{n}^{T}\bigr{)}-2\mathit{A}\nabla\cdot(\bm{n}\bm{n}),$ (6)
up to $\mathcal{O}(\gamma^{\prime},\mu_{i}^{\prime})$. This describes an isotropic Stokes flow with active and elastic driving forces [66].
In the examples presented in this letter, the fluid flow is induced by activity alone; thus, the velocity scale is set by the activity strength, $U=|\alpha|L/\mu$, and the assumption of a small Ericksen number is equivalent to assuming weak activity, $\gamma^{\prime}\mathrm{Er}=|\alpha|\gamma L^{2}/(\mu K)\ll 1$. Furthermore, in 2D, the assumption of a small viscosity ratio $\mu_{2}^{\prime}$ is unnecessary since it only corresponds to an anisotropy out-of-the-plane. Instead, $\mu_{2}$ can be absorbed into the isotropic viscosity by taking $\mu\mapsto\mu-\mu_{2}/2$.
@@
In two dimensions, the director field can be described by a director angle $\theta(x,y)\in[0,\pi)$ such that $\bm{n}=(\cos\theta,\sin\theta)$. Additionally, the conservation of mass (2b) implies that a streamfunction $\psi(x,y)$ can be introduced such that $\bm{u}=(\partial_{y}\psi,-\partial_{x}\psi)$. In this case, the equilibrium equation (5) reduces to Laplace’s equation, $\nabla^{2}\theta=0$, while the momentum equation (6) is equivalent to a Poisson’s equation for the scalar vorticity, $\omega=-\nabla^{2}\psi$. Here, we absorb the isotropic parts of the stress tensors into the pressure by taking $p\mapsto p+\mathit{A}-|\nabla\theta|^{2}/(2\mathrm{Er})$.
At any boundaries, the fluid velocity is assumed to vanish, while the director relaxes to some preferential angle $\phi$ with some anchoring strength $W$ (i.e. Rapini–Papoular anchoring [67]):
$\partial_{x}\psi=\partial_{y}\psi=0\hskip 6.00006pt\text{and}\hskip 6.00006pt\partial_{\nu}\theta=(w/2)\sin\left[2(\theta-\phi)\right],$ (7a)
for the normal derivative pointing into the LC $\partial_{\nu}=\bm{\hat{\nu}}\cdot\nabla$ and the dimensionless anchoring strength $w\coloneqq WL/K.$
Consider the complex position $z=x+\mathrm{i}y$ and its complex conjugate $\bar{z}=x-\mathrm{i}y$. Using the Wirtinger derivatives, $\partial_{x}=\partial_{z}+\partial_{\bar{z}}$ and $\partial_{y}=\mathrm{i}\partial_{z}-\mathrm{i}\partial_{\bar{z}}$, one can make a change of variables from $(x,y)$ to $(z,\bar{z})$ [68]. In these complex coordinates, the system takes the form
@@
for some locally-holomorphic functions $f$ and $g$ [68]. (Throughout, we shall assume the LC does not contain any topological defects, and so, $f$ is non-zero and singularity free inside the LC.) Inserting the imaginary part of (9b) into $\partial_{z\bar{z}}\psi=-\omega/4$ and integrating yields
@@
for another locally-holomorphic function $h$. The first two terms in (10) correspond to the well-known form of a biharmonic function with Goursat functions $h$ and $g$ [69]. The last term is a particular solution that accounts for the active forcing.
Analytical solutions for the director and velocity fields are thus obtained, provided that functions $f$, $g$, and $h$, which are locally holomorphic within the LC and satisfy the boundary conditions in (7), may be found. The problem for $f$ derived here is equivalent to the potential problem explored in Refs. [70, 71] for a passive LC. It is important to note that the solutions (9)–(10) are not invariant under a conformal mapping since they contain a mixture of holomorphic (function of $z$) and antiholomorphic (function of $\bar{z}$) terms.
As an example, consider a cylinder of radius $L$ of infinite extent immersed in a weakly-active nematic LC. In the far-field, we assume the LC is aligned horizontally in a direction orthogonal to the cylinder’s long axis and the fluid velocity vanishes. Additionally, the LC is assumed to be tangentially anchored to the cylinder with a finite strength, i.e. (7b) with $\phi=\arg(\mathrm{i}z)\mod\pi$ on $|z|=1$. Without the body, the system sits at an equilibrium, with $\bm{n}=\bm{\hat{x}}$ and $\bm{u}=\bm{0}$, and the active stress is absorbed by the pressure.
First, we determine the LC director angle $\theta=-\arg f^{\prime}$. Here, the potential $f(z)$ must be locally holomorphic in $1<|z|<\infty$ and satisfy the finite anchoring condition on $|z|=1$, and $f(z)\sim z$ as $|z|\to\infty$. The solution to this problem that minimizes the Frank elastic energy is, for any $w>0$,
@@
where $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$ is an effective radius [70]. The corresponding director lines (which lay parallel to $\bm{n}$) are plotted in Fig. 1(a) for $w=10$.
Next, with (11), we determine the activity-induced fluid flow. Here, the Goursat functions $h(z)$ and $g(z)$ must be locally holomorphic in $1<|z|<\infty$ and yield a fluid velocity that vanishes on $|z|=1$ and as $|z|\to\infty$, ultimately yielding the streamfunction
$\psi=\frac{\mathit{A}\rho^{2}}{8}\operatorname{Im}\biggl{[}\frac{(|z|^{2}-1)^{2}}{z^{2}(z^{2}-\rho^{2})}+\frac{1-|z|^{2}+2\log|z|}{(z^{2}-\rho^{2})/(2\rho^{2})}\biggr{]},$ (12)
@@
see [66]. The corresponding velocity field is shown in Fig. 1(b,c) for $w=1$ and $w=10$, with $\mathit{A}=-1$ (i.e. extensile activity). A prominent feature is a plume extending symmetrically to both the left and right of the cylinder. Far from the body, the velocity field behaves as
$u+\mathrm{i}v\sim\bigl{[}\rho^{2}\cos(2\alpha)-\cos(4\alpha)\bigr{]}\mathit{A}\rho^{2}e^{\mathrm{i}\alpha}/(2|z|),$ (13)
@@
as $|z|\to\infty$, for $\alpha=\arg z$; the plume angle thus decreases from $\pi/4$ to 00 as the anchoring strength $w$ is increased from 00 to $\infty$, while the plume speed is largest at $w=8/3$. The velocity field does not appear as a dipole or stresslet in the far-field — the mere presence of the body initiates additional flow far away due to the distributed active stress, leading to bulk flow release. This also demonstrates that a body with non-trivial anchoring conditions is sufficient to induce a large scale flow through its presence alone.
Both the elasticity associated with passive LCs and the flow induced by activity generate stresses that act on any boundaries. Here, the passive LC introduces a surface traction $\bm{t}^{e}=\bm{\hat{\nu}}\cdot\hat{\bm{\sigma}}^{e}+\partial_{s}\bm{t}^{s}$, where $\hat{\bm{\sigma}}^{e}=(1/2)\lvert\nabla\theta\rvert^{2}\bm{\mathrm{I}}-\nabla\theta\nabla\theta$ is the (traceless) elastic stress tensor and $\bm{t}^{s}$ is a surface stress vector caused by the finite surface-anchoring (7b) [66]. The activity generates an additional surface traction $\bm{t}^{a}=\mathrm{Er}\,\bm{\hat{\nu}}\cdot(\hat{\bm{\sigma}}^{v}+\mathit{A}\hat{\bm{\sigma}}^{a}-p\bm{\mathrm{I}})$ for the (traceless) viscous and active stress tensors $\hat{\bm{\sigma}}^{v}=\nabla\bm{u}+\nabla\bm{u}^{T}$ and $\hat{\bm{\sigma}}^{a}=2\bm{n}\bm{n}-\bm{\mathrm{I}}$, respectively.

### v3

@@ Abstract
A body immersed in a nematic liquid crystal disturbs the fluid’s preferred molecular configuration and increases its stored elastic energy. In an active nematic, the fluid components also generate a stress in the bulk fluid. By introducing either an immersed body or boundary, a large scale flow can be triggered due to anchoring boundary conditions alone — a global pressure built by active stresses at equilibrium is instantly released everywhere. The fluid then imposes viscous, elastic, and active stresses on such surfaces which, if compliant, may result in a surface deformation. We study the deformations and stresses of a linearly elastic body placed in an active nematic in two dimensions. Using complex variables techniques, exact expressions for the fluid flow, director field, surface tractions, and body deformation are derived. Qualitative differences between elastic and active stress-driven deformations are identified, depending on an active Ericksen number, anchoring conditions, and body material properties, thereby suggesting a new method for measuring mechanical stresses in active anisotropic environments. Flow profiles, external confinement, and anchoring-induced stirring are also addressed.
I Introduction
@@
The eukaryotic cell is host to a wide variety of deformable organelles, from the nucleus to mitochondria [1, 2, 3, 4, 5, 6]. These organelles reside in complex environments which can be anisotropic [7, 8], and experience mechanical stresses due to actively driven cytoplasmic flows [9, 10, 11, 12, 13, 14, 15]. The cell membrane itself is also deformable, and its morphology is commonly influenced by internal processes. For instance, active anisotropic stresses are of functional importance in cytoskeletal remodeling of the nucleus and chromatin configurations [16, 17, 18, 19, 20], in the metaphase spindle [21, 22, 23], and in numerous other aspects of tissue morphogenesis [24, 25, 26]. On a larger length scale, active systems like bacterial swarms and biofilms also exhibit anisotropy, among other complex rheological properties [27, 28, 29, 30, 31].
@@
Even when free of bodies, actively-driven anisotropic flows exhibit very rich and complex dynamics [32, 33, 34, 35, 36]. Stability [37, 34, 38, 39, 40, 41, 42, 43], topology [44], pattern formation [45, 46, 47], and mixing [48, 49] have seen substantial theoretical treatment. Confinement introduces an additional length scale, which can affect the emergent spatial structure [50, 51]. Strong active nematic stresses can also change the shape of deformable confining surfaces [52, 53, 54, 55, 56, 57, 58, 59], and can play a prominent role in directing the motility of cells [60, 61] and active droplets [62, 63, 64, 65].
@@
When an active nematic contains an immersed body, or inclusion, additional elastic stresses are immediately introduced. In passive systems, these stresses result in intricate and beautiful textures [66, 67, 68], and can be used to measure cell material properties [69, 70]. In active systems, fluid elasticity and anisotropy can redirect the emergent flows, leading to complex body dynamics. Recent examples include an investigation of the biased rotation of gear-like bodies [71, 72], complex droplet trajectories [73, 74], and transitions from fixed point to limit cycle to chaotic dynamics of a circular disk [75]. How body deformability affects and is affected by such an active environment remains largely unexplored.
In this paper, we study an active nematic fluid based on Ericksen–Leslie theory [76, 77, 78, 79, 36] with a soft internal or external boundary in two dimensions. Complex variable techniques yield exact analytical representations for the director field, velocity field, and boundary deformation. Elastic deformations of bounding surfaces are known to relieve the elastic energy stored in the bulk LC phase [80]. Here, we show that active stresses can compete and eventually dominate this balance, resulting in distinct geometric signals. In addition, we show that a spontaneous stirring flow with a unique plume structure may be induced with a change in surface anchoring boundary conditions alone, due to a global pressure release.
II Active Nematic Flows
@@
Active nematic liquid crystals (LCs) can be described by their locally-averaged molecular orientation (i.e. the director field) $\bm{n}(\bm{x},t)$ and fluid velocity $\bm{u}(\bm{x},t)$, with spatial position $\bm{x}$, time $t$, and $|\bm{n}|=1$. The director field, $\bm{n}$, evolves due to gradients in the velocity field, with strain-rate tensor $E_{ij}=(\partial_{i}u_{j}+\partial_{j}u_{i})/2$ and vorticity tensor $\Omega_{ij}=(\partial_{i}u_{j}-\partial_{j}u_{i})/2$, and a relaxation towards equilibrium under the action of the molecular field, $\bm{H}=-\delta\mathcal{F}/\delta\bm{n}$, where $\mathcal{F}$ is the free energy density. In the one-constant approximation, $\mathcal{F}=K\lVert\nabla\bm{n}\rVert^{2}/2$ and so $\bm{H}=K\nabla^{2}\bm{n}$, for the single (Frank) elastic constant $K$ [77]. Together,
$(\partial_{t}+\bm{u}\cdot\nabla)\bm{n}=\bm{n}\cdot\bm{\Omega}+\lambda\bm{n}\cdot\bm{\mathrm{E}}\cdot(\bm{\mathrm{I}}-\bm{n}\bm{n})+\bm{h}/\gamma,$ (1)
@@
with $\gamma$ the rotational viscosity, $\lambda$ the tumbling/reactive parameter, and $\bm{h}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\bm{H}$ the transverse part of the molecular field, where $\bm{n}\bm{n}$ is a dyadic product [76, 77, 78, 80]. Momentum and mass conservation, under the assumptions of incompressibility and zero Reynolds number flow [^1] The Reynolds number is defined as $\operatorname{Re}=\rho|\alpha|L^{2}/\mu^{2}$, with $\rho$ the fluid density, and other terms defined above., are given by
@@
respectively, for $\nabla\cdot\bm{\sigma}\coloneqq\partial_{i}\sigma_{ij}$, pressure $p$, and elastic, viscous, and active stress tensors
$\begin{split}\bm{\sigma}^{e}&=\frac{K}{2}\lVert\nabla\bm{n}\rVert^{2}\mathbf{I}-K\nabla\bm{n}\cdot\nabla\bm{n}^{T}\\ &\quad-\frac{1}{2}(\bm{n}\bm{h}-\bm{h}\bm{n})-\frac{\lambda}{2}(\bm{n}\bm{h}+\bm{h}\bm{n}),\end{split}$ (3a)
@@
respectively. Here, $\mu$ is the solvent viscosity, $\mu_{1}$ and $\mu_{2}$ are anistropic viscosities, and $\alpha$ is the activity strength, which can describe both extensile ($\alpha<0$) and contractile ($\alpha>0$) activity [79, 81, 36] (see [82]). Note the convention taken for the divergence, which is important since $\bm{\sigma}^{e}$ is not symmetric.
The system is made dimensionless by scaling upon an intrinsic length scale $L$ (e.g. the size of an immersed body), velocity scale $U$ (to be defined), and stress scale $K/L^{2}$. The system is then characterized by the dimensionless Ericksen number, activity strength, and viscosity ratios [82]:
@@
For small rotational Ericksen numbers, $\gamma^{\prime}\mathrm{Er}\ll 1$, the fluid flow does not affect the director field, and at leading order from (1) we have
$\bm{h}=(\mathbf{I}-\bm{n}\bm{n})\cdot\nabla^{2}\bm{n}=\bm{0}.$ (5)
Further, assuming small viscosity ratios, $\gamma^{\prime},\mu_{i}^{\prime}\ll 1$, Eq. (2a) yields
$-\nabla p+\nabla^{2}\bm{u}+2\mathit{A}\nabla\cdot(\bm{n}\bm{n})=\bm{0}.$ (6)
In the examples presented in this paper, the fluid flow is induced by activity alone; thus, the velocity scale is set by the activity strength, $U=|\alpha|L/\mu$, and the assumption of a small Ericksen number is equivalent to assuming weak activity, $\gamma^{\prime}\mathrm{Er}=|\alpha|\gamma L^{2}/(\mu K)\ll 1$.  [^2] In 2D incompressible flows, a general $\mu_{2}$ can actually be absorbed into the isotropic viscosity by taking $\mu\mapsto\mu-\mu_{2}/2$.
@@
II.1 Two-dimensional active nematics
In 2D, the director field can be described by an angle field $\theta(x,y)\in[0,\pi)$ such that $\bm{n}=(\cos\theta,\sin\theta)$, and a streamfunction $\psi(x,y)$ can be introduced such that $\bm{u}=(\partial_{y}\psi,-\partial_{x}\psi)$. The director field, governed by $\bm{h}=\bm{0}$, reduces to Laplace’s equation for $\theta$, while the momentum equation (6) is equivalent to Poisson’s equation for the scalar vorticity, $\omega=-\nabla^{2}\psi$:
$\qquad\nabla^{2}\theta=0,$ (7a)
$\nabla^{2}\omega=\mathit{A}\left[2\partial_{yx}\cos(2\theta)+(\partial_{yy}-\partial_{xx})\sin(2\theta)\right].$ (7b)
Here, we absorb the isotropic parts of the stress tensors into the pressure by taking $p\mapsto p+\mathit{A}$. The leading-order, dimensionless, traceless stress tensors are denoted with hats:
$\bm{\hat{\sigma}}^{e}$ $=\frac{1}{2}|\nabla\theta|^{2}\mathbf{I}-\nabla\theta\nabla\theta,$ (8a)
$\bm{\hat{\sigma}}^{v}$ $=\nabla\bm{u}+\nabla\bm{u}^{T},$ (8b)
$\bm{\hat{\sigma}}^{a}$ $=2\bm{n}\bm{n}-\mathbf{I},$ (8c)
giving the dimensionless (bulk) stress tensor $\bm{\sigma}=\bm{\hat{\sigma}}^{e}+\mathrm{Er}\left(\bm{\hat{\sigma}}^{v}+A\bm{\hat{\sigma}}^{a}-p\mathbf{I}\right)$.
At any boundaries, we assume the fluid velocity vanishes and the director field relaxes locally towards a preferred angle $\theta_{0}(s)$ due to surface anchoring of strength $W$ (i.e. Rapini–Papoular anchoring [83]):
$\partial_{s}\psi=\partial_{\nu}\psi=0\quad\text{and}\quad\partial_{\nu}\theta=\frac{w}{2}\sin\left[2(\theta-\theta_{0})\right],$ (9a)
for the anticlockwise arclength derivative $\partial_{s}=\bm{\hat{s}}\cdot\nabla$, normal derivative pointing into the LC $\partial_{\nu}=\bm{\hat{\nu}}\cdot\nabla$, and the dimensionless anchoring strength $w\coloneqq WL/K.$
With $z=x+\mathrm{i}y$ and $\bar{z}=x-\mathrm{i}y$ its complex conjugate, the system may be written in complex variables as
@@ II.1 Two-dimensional active nematics
for some locally-holomorphic functions $f$ and $g$ [84]. We assume the LC does not contain any topological defects, so $f$ is non-zero and singularity free inside the LC, though this assumption can be relaxed [85, 86]. Inserting the imaginary part of (11b) into $\partial_{z\bar{z}}\psi=-\omega/4$ and integrating yields
@@
for another locally-holomorphic function $h$. The first two terms in (12) correspond to the well-known form of a biharmonic function with Goursat functions $h$ and $g$ [87], while the last term accounts for the active forcing.
Analytical solutions for the director and velocity fields are obtained, provided that functions $f$, $g$, and $h$, which are locally holomorphic within the LC and satisfy the boundary conditions in (9), may be found. The problem for $f$ derived here is equivalent to the potential problem explored in Refs. [88, 89] for a passive LC, but $g$ and $h$ depend on activity.
II.2 Anchoring-controlled flows
As an example, consider a cylinder of radius $L$ of infinite extent immersed in a weakly-active nematic LC. In the far-field, we assume the LC is aligned horizontally in a direction orthogonal to the cylinder’s long axis and the fluid velocity vanishes. The LC is assumed to be tangentially anchored to the body with finite strength, i.e. (9b) with $\theta_{0}=\arg(\mathrm{i}z)\mod\pi$ on $|z|=1$. Without the body, the system sits at an equilibrium, with $\bm{n}=\bm{\hat{x}}$ and $\bm{u}=\bm{0}$, and the active stress is absorbed by the pressure.
To find the activity-induced flow, we first determine the LC director angle $\theta=-\arg f^{\prime}$. Here, the potential $f(z)$ must be locally holomorphic in $1<|z|<\infty$ and satisfy the finite anchoring condition on $|z|=1$, and $f(z)\sim z$ as $|z|\to\infty$. The solution to this problem is
@@ II.2 Anchoring-controlled flows
where $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$ is an effective radius [88]. The corresponding director lines (which lay parallel to $\bm{n}$) are plotted in Fig. 1(a) for $w=10$. Meanwhile, the Goursat functions $h(z)$ and $g(z)$ must be locally holomorphic in $1<|z|<\infty$ and yield a fluid velocity that vanishes on $|z|=1$ and as $|z|\to\infty$. This ultimately yields the streamfunction (see [82]):
$\psi=\frac{\mathit{A}\rho^{2}}{8}\operatorname{Im}\biggl{[}\frac{(|z|^{2}-1)^{2}}{z^{2}(z^{2}-\rho^{2})}+\frac{1-|z|^{2}+2\log|z|}{(z^{2}-\rho^{2})/(2\rho^{2})}\biggr{]}.$ (14)
The corresponding flow is shown in Fig. 1(b,c) for $w=1$ and $w=10$, with $\mathit{A}=-1$ (i.e. extensile activity).
@@
A prominent feature of the induced flow is a plume extending symmetrically to both the left and right of the cylinder (see Fig. 1(b,c)). Far from the body,
$u+\mathrm{i}v=-2\mathrm{i}\partial_{\bar{z}}\psi\sim\bigl{[}\rho^{2}\cos(2\vartheta)-\cos(4\vartheta)\bigr{]}\frac{\mathit{A}\rho^{2}e^{\mathrm{i}\vartheta}}{2|z|},$ (15)
@@
as $|z|\to\infty$, for $\vartheta=\arg z$. This far-field flow behavior is critically dependent on the effective radius $\rho(w)$, and hence the anchoring strength $w$. The plume angle, thus, decreases from $\pi/4$ to 00 as the anchoring strength $w$ is increased from 00 to $\infty$, while the plume speed is largest at $w=8/3$. The velocity field does not appear as a dipole or stresslet in the far-field, but the sum of a dipole and a quadrupole — the mere presence of the body initiates additional flow far away due to the distributed active stress, leading to bulk flow release. This also demonstrates that a body with non-trivial anchoring conditions is sufficient to induce a large scale flow through its presence alone.
III Surface tractions and soft boundaries
Stresses due to passive LC elasticity, actively-driven viscous flow, and direct activity all contribute to surface forcing. The LC introduces a surface traction $\bm{t}^{e}=\bm{\hat{\nu}}\cdot\bm{\hat{\sigma}}^{e}+\partial_{s}\bm{t}^{s}$, where $\bm{\hat{\sigma}}^{e}$ is the Ericksen elastic stress tensor given in (8) and
$\bm{t}^{s}=\frac{w}{2}\sin^{2}(\theta-\theta_{0})\bm{\hat{s}}+\frac{w}{2}\sin[2(\theta-\theta_{0})]\bm{\hat{\nu}},$ (16)
is a surface stress vector associated with finite surface-anchoring [80]. Activity, meanwhile, produces a surface traction $\bm{t}^{a}=\mathrm{Er}\,\bm{\hat{\nu}}\cdot(\bm{\hat{\sigma}}^{v}+\mathit{A}\bm{\hat{\sigma}}^{a}-p\bm{\mathrm{I}})$ for the (traceless) viscous and active stress tensors, $\bm{\hat{\sigma}}^{v}$ and $\bm{\hat{\sigma}}^{a}$ given in (8).